A cylinder containing CH4, C2H6 and N2 has to be prepared containing a mole ratio of CH4 to C2H6 of 0,5 to 1. What is the mol ratio of N2 and CH4 and the mol ratio of N2 and C2H6 in the mixture?

Available to prepare the mixture are:

1. a cylinder containing a mixture of 65% N2 and 35% CH4

2. a cylinder containing a mixture of 70% N2 and 30% C2H6


1
Expert's answer
2021-03-22T06:28:28-0400

1. % by mass of N2= 65

% by mass of CH4= 35

Divide by their molar masses

Mole of N2= 65/28= 2.32

Mole of CH4= 35/16= 2.19

Divide by the smaller

N2= 2.32/2.19= 1.06

CH4= 2.19/2.19= 1

The mole ratio of N2 to CH4 is approximately 1:1


2. % by mass of N2= 70

% by mass of C2H6= 30

Divide by their molar masses

Mole of N2= 70/28= 2.5

Mole of C2H6= 30/28= 1.07

Divide by the smaller

N2= 2.5/1.07= 2.33

C2H6= 1.07/1.07=1

Multiply by a factor of 3

N2= 2.33 x 3= 7

C2H6= 1x3= 3

Mole ratio of N2 to C2H6 is 7:3